# Bridging the Synthesizability Gap in Perovskites by Combining Computations, Literature Data, and PU Learning

### v1

@@ 4 Future Outlook
Given the expense of running DFT simulations at GGA accuracy or higher-level functionals, especially using large supercells to simulate alloys or defects, it is vital to implement ML-driven acceleration. For instance, DFT geometry optimization can be augmented with graph neural network (GNN)-based interatomic potentials [43, 76], such as the models recently developed in our group [44]. GNN+DFT will be used to perform efficient structural relaxation to obtain ground state configurations. Interatomic potentials can further aid in navigating the defect landscape and performing perturbation-based calculations. The final stage would involve performing rigorous high-accuracy DFT computations exclusively for the curated subset of best materials. Additionally, these compounds will be recommended to experimental collaborators for synthesis and characterization; any new experimental data will serve to validate the predicted synthesizability and computed properties. This multi-stage computational screening approach is designed to dramatically reduce the number of required calculations, thereby accelerating both the simulations and the discovery of novel perovskite materials.
@@
  • Extension to other DFT functionals, including hybrid HSE06 calculations with and without spin-orbit coupling [77] to improve electronic band edges and gaps; this will lead to multi-fidelity models combining GGA and HSE06 data.
  • An expanded chemical space which includes a series of other possible A/B/B’ cations, such as 3d, 4d, and 5d transition metals, and a variety of other organic molecular cations [78, 79, 80].
@@
  • Approaches other than PU learning can be explored to make synthesis probability predictions [48]; additionally, the synthesizability framework can be integrated with structure-based GNN models to eliminate the need for hand-crafted descriptors, especially as the dataset swells in size.
@@ 7 Data Availability
Data and code can also be accessed using our nanoHUB tool [81].

### v2

@@ 4 Future Outlook
Given the expense of running DFT simulations at GGA accuracy or higher-level functionals, especially using large supercells to simulate alloys or defects, it is vital to implement ML-driven acceleration. For instance, DFT geometry optimization can be augmented with graph neural network (GNN)-based interatomic potentials [m3gnet, mace-off], such as the models recently developed in our group [perovsiap]. GNN+DFT will be used to perform efficient structural relaxation to obtain ground state configurations. Interatomic potentials can further aid in navigating the defect landscape and performing perturbation-based calculations. To accelerate our DFT workflows, we aim to develop a unified model by employing a fine-tuning strategy analogous to that of the current Perovs-IAP framework. This new model will be trained on a comprehensive dataset encompassing all types of perovskite structures. The final stage would involve performing rigorous high-accuracy DFT computations exclusively for the curated subset of best materials. Additionally, these compounds will be recommended to experimental collaborators for synthesis and characterization; any new experimental data will serve to validate the predicted synthesizability and computed properties. This multi-stage computational screening approach is designed to dramatically reduce the number of required calculations, thereby accelerating both the simulations and the discovery of novel perovskite materials.
@@
  • Extraction of more experimentally known compounds from the literature will continue, especially using new publications that emerge every day.
  • Extension to other DFT functionals, including hybrid HSE06 calculations with and without spin-orbit coupling [soc] to improve electronic band edges and gaps; this will lead to multi-fidelity models combining GGA and HSE06 data.
  • An expanded chemical space which includes a series of other possible A/B/B’ cations, such as 3d, 4d, and 5d transition metals, and a variety of other organic molecular cations [mol-1, mol-2, dmetal].
@@
  • Approaches other than PU learning can be explored to make synthesis probability predictions [syncotrain]; additionally, the synthesizability framework can be integrated with structure-based GNN models to eliminate the need for hand-crafted descriptors, especially as the dataset swells in size.
@@
  • Finally and very importantly, it is vital that our predictions lead to experimental validation and discovery of new perovskites. Efforts are currently underway for synthesizing some of the compounds predicted to have the highest synthesis probabilities, and we welcome interest from the community in taking our predictions to the laboratory.
@@ 7 Data Availability
Data and code can also be accessed using our nanoHUB tool [nh-tool].